OpenStudy (anonymous):

In a spherical electromagnetic wave, what is the configuration of the electric and magnetic fields? I was thinking perhaps the electric field could orient itself around a hemisphere, but then following the pattern would cause destructive interference at the poles, so that can't work. The Huygens–Fresnel principle is based on the notion of spherical waves, but how does a spherical wave work in the context of electromagnetic wave propagation?

OpenStudy (anonymous):

Huygens was talking about wavefronts and their propagation

OpenStudy (anonymous):

"In 1678, Huygens proposed that every point to which a luminous disturbance reaches becomes a source of a spherical wave; the sum of these secondary waves determines the form of the wave at any subsequent time." Source: http://en.wikipedia.org/wiki/Huygens%27_principle Either way, Huygen's principle really isn't my question....it's just what triggered it. I'm interested in the notion of the spherical waves.

OpenStudy (vincent-lyon.fr):

What you say is true and very insightful ! An electromagnetic spherical wave, emitted by an oscillating dipole is not isotropic. There is actually no energy emitted along the axis of the dipole. Mean Poynting vector varies as sin²θ, so energy is maximum in the equatorial plane of the dipole. This is the reason why light diffused by air molecules in the sky is partly polarised. That spherical wave, as soon as you are far enough form the dipole has the local structure of a plane wave (E and B transverse and perpendicular to each other ; \(\vec B=\Large \frac {\vec u}{c}\normalsize \times \vec E\) ) \(\vec E\) is along the \(\vec e_\theta\) unit vector \(\vec B\) is along the \(\vec e_\phi\) unit vector both becoming zero as θ approaches 0 or π. This is somehow linked to the Hairy Ball Theorem. See: http://www.youtube.com/watch?v=B4UGZEjG02s On the contrary, a sound wave, being longitudinal, could be spherical AND isotropic. You need a pulsing sphere to produce it. but you cannot produce an isotropic spherical transverse wave. In the case of light, a real source is made up of zillions of non coherent atoms and the light wave produced will be isotropic and non polarised. (This effect is somehow similar to the one you described in you question about random polarisation that was solved by @Jemurray3). Hope this helps!

OpenStudy (anonymous):

Thank you! That was very helpful!!